# Deneb: A Hallucination-Robust Automatic Evaluation Metric for Image Captioning

### v1

@@ 1 Introduction
Therefore, we propose Deneb [^1] In the field of general text generation metrics, names of celestial bodies are often used as names for metrics, such as METEOR[6] and COMET[44]. Following this convention, we have named our metric and dataset after a celestial body., a supervised automatic evaluation metric specifically robust against hallucinations. Fig. 1 shows an overview of Deneb. Unlike pseudo-multifaceted metrics, Deneb is a multifaceted metric and thus can effectively compare multifaceted descriptions of an image against a candidate caption. This capability is achieved by the Sim-Vec Transformer, a novel architecture designed for the efficient processing of multiple reference captions. Therefore, we construct the Nebula dataset by extending the Polaris dataset, increasing the visual diversity by a factor of three. Our dataset offers more effective data for training supervised metrics, particularly those capable of handling a wide range of cases, including hallucinated captions.
@@
    We propose Sim-Vec Extraction (SVE), which utilizes a Hadamard product and element-wise differences to extract features beneficial for automatic evaluation.
@@
    We achieve state-of-the-art performance among LLM-free metrics on FOIL, Composite, Flickr8K-Expert, Flickr8K-CF, PASCAL-50S, and the Nebula dataset.
@@ 2 Related Work
A notable metric developed to evaluate this issue is CHAIR [45], which assesses hallucinations by calculating the proportion of generated words that accurately reflect objects in the input image. However, CHAIR’s approach is rule-based and operates within a closed vocabulary, limiting its applicability and generalizability. The development of CHAIR underscores the need for more comprehensive benchmarks to evaluate the robustness of automatic evaluation metrics against hallucinations.
In response to the challenge of hallucinations, [48] introduced the FOIL benchmark to evaluate the robustness of metrics against hallucinations. The FOIL dataset, derived from the COCO dataset [33], consists of approximately 200,000 image-caption pairs, featuring a mix of both correct and hallucinated captions. Hallucinated captions are generated by altering a single word in a correct caption to create a similar but inaccurate version — for instance, replacing “motorcycle” with “bicycle”.
Moreover, the study in [19] presented a widely-accepted methodology for using the FOIL dataset to evaluate metrics. As detailed in [19], this methodology entails the following steps: For each of the 32K test images in the dataset, a pair comprising a hallucinated caption and its correct counterpart is sampled. The robustness of each evaluation metric against hallucinations is then evaluated based on its capability to consistently assign higher scores to the correct captions than to their hallucinated versions.
@@
Standard datasets for evaluating image captioning metrics include FOIL [48], Flickr8K-Expert [20], Flickr8K-CF, Composite [1], Polaris [55], and PASCAL-50S [4]. The FOIL dataset, specifically designed to assess metric robustness against hallucination, is derived from the COCO dataset and comprises approximately 200,000 image-caption pairs with a mix of correct and hallucinated captions. Flickr8K-Expert, Flickr8K-CF, Composite, and Polaris each consist of human Likert-scale judgments at the level of each image-caption pair. Specifically, Flickr 8K-Expert comprises 17,000 human judgments across 5,664 images, with captions rated on a four-point scale. Flickr8K-CF offers 145,000 human judgments collected from CrowdFlower, covering over 48,000 image-caption pairs. The Composite dataset, sourcing image-caption pairs from COCO, Flickr8K [20], and Flickr30K [60], contains approximately 12,000 human judgments. In response to concerns regarding the lack of diversity in such datasets, the Polaris dataset [55] was introduced, and it is currently the largest dataset in the field. The Polaris dataset is tailored for the training and evaluation of metrics and consists of a diverse collection of captions, sourced from ten modern image captioning models and accompanied by approximately 130,000 human judgments. However, Polaris faces the challenge of a limited variety of images compared with the total number of samples, leading to concerns about imbalance in the dataset.
@@ 4.1 Experimental Setup
The development of supervised metrics for image captioning requires a large-scale, diverse dataset, but there are limited available datasets for effectively training these models. Moreover, the largest dataset in this field[55] has a notable issue: a significant imbalance, characterized by a limited variety of images relative to the number of captions. Specifically, there is a discrepancy where the number of images is approximately only one-tenth of the total number of samples. This imbalance could potentially lead to suboptimal evaluation of various types of images. To mitigate this imbalance, we constructed the Nebula dataset by extending the Polaris dataset to have approximately three times the number of images.
Following the standard procedure, human judgments were adopted on a five-point scale to assess the appropriateness of a candidate for a given image and references. The annotation process was carried out through a crowdsourcing service. Following previous studies [27, 54, 55], we instructed the annotators to assess the quality of the candidates from the perspectives of fluency, relevance, and descriptiveness. To ensure the reliability of our data, we excluded data from evaluators who exhibited suspicious behavior, such as extremely short response times or consistently providing identical values. In addition, the human judgments, given on a five-point scale, were normalized to the range $[0,1]$.
The Nebula dataset comprises 32,978 images and 32,978 human judgments collected from 805 annotators, and contains approximately three times more images than the Polaris dataset. The total number of references is 183,472, with a vocabulary size of 32,870, a total word count of 1,945,956, and an average sentence length of 10.61 words. The total number of candidates is 32,978, with a vocabulary size of 3,695, a total word count of 288,922, and an average sentence length of 8.76 words. All sentences are in English. For a comprehensive description of the dataset, please refer to the Appendix.
Implementation Details.
Our model had approximately 133 million trainable parameters. We trained our model on a Tesla A100 GPU and measured the inference time on the GeForce RTX 3090 with 24 GB of memory and an Intel Core i9 12900K with 64 GB of memory. The training phase was completed in approximately 2 hours, and the inference time per sample on GeForce RTX 3090 was approximately 22 ms. Further details are explained in Appendix.
@@ 4.2 Comparison to State-of-the-Art
Figure 3: Qualitative results on the Nebula dataset. Panels (a) and (b) feature images from successful cases, and panel (c) depicts a failure case.
@@ 4.4 Ablation Studies
Appendix 0.A Implementation Details
@@
Appendix 0.B Nebula Dataset
@@
Appendix 0.C Additional Qualitative Analysis
@@
Appendix 0.D Error Analysis

### v2

@@ 1 Introduction
Therefore, we propose Deneb [^1] In the field of general text generation metrics, names of celestial bodies are often used as names for metrics, such as METEOR[6] and COMET[44]. Following this convention, we have named our metric and dataset after a celestial body., a supervised automatic evaluation metric specifically robust against hallucinations. Fig. 1 shows an overview of Deneb. Unlike pseudo-multifaceted metrics, Deneb is a multifaceted metric and thus can effectively compare multifaceted descriptions of an image against a candidate caption. This capability is achieved by the Sim-Vec Transformer, a novel architecture designed for the efficient processing of multiple reference captions.
Furthermore, for training our proposed metric, we construct the Nebula dataset by extending the Polaris dataset, increasing the visual diversity by a factor of three. Our dataset offers more effective data for training supervised metrics, particularly those capable of handling a wide range of cases, including hallucinated captions.
@@
    We propose Sim-Vec Extraction (SVE), which utilizes a Hadamard product and element-wise differences to extract features beneficial for evaluation.
@@
    We achieve state-of-the-art performance among LLM-free metrics on FOIL, Composite, Flickr8K-Expert, Flickr8K-CF, PASCAL-50S, and Nebula.
@@ 2 Related Work
A notable metric developed to evaluate this issue is CHAIR [45], which assesses hallucinations by calculating the proportion of generated words that accurately reflect objects in the input image. However, CHAIR’s approach is rule-based and operates within a closed vocabulary, limiting its applicability and generalizability. The development of CHAIR underscores the need for more comprehensive benchmarks to evaluate the robustness against hallucinations.
In response to the challenge of hallucinations, Shekhar et al. [48] introduced the FOIL benchmark to evaluate the robustness of metrics against hallucinations. The FOIL dataset, derived from the COCO dataset [33], consists of approximately 200,000 image-caption pairs, featuring a mix of both correct and hallucinated captions. Hallucinated captions are generated by altering a single word in a correct caption to create a similar but inaccurate version — for instance, replacing “motorcycle” with “bicycle”. Moreover, Hessel et al. [19] presented a widely-accepted methodology for using the FOIL dataset to evaluate metrics. As detailed in [19], this methodology entails the following steps: For each of the 32K test images in the dataset, a pair comprising a hallucinated caption and its correct counterpart is sampled. The robustness of each evaluation metric against hallucinations is then evaluated based on its capability to consistently assign higher scores to the correct captions than to their hallucinated versions.
@@
Standard datasets for evaluating image captioning metrics include FOIL [48], Flickr8K-Expert [20], Flickr8K-CF, Composite [1], Polaris [55], and PASCAL-50S [4]. In response to concerns about the lack of diversity in the Flickr8K-Expert, Flickr8K-CF, and Composite datasets, Wada et al. [55] introduced the Polaris dataset, which is currently the largest in the field. However, as we argue later, Polaris faces the challenge of a limited variety of images compared with the total number of samples, leading to concerns about imbalance in the dataset.
@@ 4.1 Experimental Setup
The development of supervised metrics for image captioning requires a large-scale, diverse dataset, but there are limited available datasets for effectively training these models. Moreover, the largest dataset in this field[55] has a notable issue: a significant imbalance, characterized by a limited variety of images relative to the number of captions. Specifically, there is a discrepancy where the number of images is approximately only one-tenth of the total number of samples. This imbalance could potentially lead to suboptimal evaluation of various types of images.
To mitigate this imbalance, we constructed the Nebula dataset [^2] The Nebula dataset can be downloaded from this link. by extending the Polaris dataset to have approximately three times the number of images. Following the standard procedure, human judgments were adopted on a five-point scale to assess the appropriateness of a candidate for a given image and references. The annotation process was carried out through a crowdsourcing service. Following previous studies [27, 54, 55], we instructed the annotators to assess the quality of the candidates from the perspectives of fluency, relevance, and descriptiveness. To ensure the reliability of our data, we excluded data from evaluators who exhibited suspicious behavior, such as extremely short response times or consistently providing identical values. In addition, the human judgments, given on a five-point scale, were normalized to the range $[0,1]$. For a comprehensive description of the dataset, please refer to the Appendix.
@@ 4.2 Comparison to State-of-the-Art
Figure 3: Qualitative results on the Nebula dataset. Panels (a) and (b) illustrate successful cases, and panel (c) depicts a failure case.
@@ 4.4 Ablation Studies
Table 3: Quantitative results of the ablation studies. The incorporation of the SVE Transformer, as indicated for Metrics (i) and (ii), emerges as the most significant factor in enhancing performance.
@@
Acknowledgements
This work was partially supported by JSPS KAKENHI Grant Number 23H03478, JST CREST, and NEDO.
Appendix 0.A Additional Related Work
The FOIL dataset, specifically designed to assess metric robustness against hallucination, is derived from the COCO dataset and comprises approximately 200,000 image-caption pairs with a mix of correct and hallucinated captions. Flickr8K-Expert [20], Flickr8K-CF, Composite[1], and Polaris[55] each consist of human Likert-scale judgments at the level of each image-caption pair. Specifically, Flickr 8K-Expert comprises 17,000 human judgments across 5,664 images, with captions rated on a four-point scale. Flickr8K-CF offers 145,000 human judgments collected from CrowdFlower, covering over 48,000 image-caption pairs. The Composite dataset, sourcing image-caption pairs from COCO[33], Flickr8K [20], and Flickr30K [60], contains approximately 12,000 human judgments. The Polaris dataset is tailored for the training and evaluation of metrics and consists of a diverse collection of captions, sourced from ten modern image captioning models and accompanied by approximately 130,000 human judgments.
Appendix 0.B Implementation Details
@@
Our model had approximately 133 million trainable parameters. We trained our model on a Tesla A100 GPU and measured the inference time on the GeForce RTX 3090 with 24 GB of memory and an Intel Core i9 12900K with 64 GB of memory. The training phase was completed in approximately 2 hours, and the inference time per sample on GeForce RTX 3090 was approximately 22 ms.
Appendix 0.C Nebula Dataset
@@
The Nebula dataset comprises 32,978 images and 32,978 human judgments collected from 805 annotators, and contains approximately three times more images than the Polaris dataset. The total number of references is 183,472, with a vocabulary size of 32,870, a total word count of 1,945,956, and an average sentence length of 10.61 words. The total number of candidates is 32,978, with a vocabulary size of 3,695, a total word count of 288,922, and an average sentence length of 8.76 words. All sentences are in English.
Appendix 0.D Additional Qualitative Analysis
@@
Appendix 0.E Error Analysis